find the equation of a straight line that passes through the midpoint of (0 , 4 ) and (6 , 0) and has a gradient of -3
Answer by Cromlix(4381)

Mid point of (0,4) and 6,0) = 6/2,4/2
= 3, 2.
Using equation y = mx + c
y = -3x + c
Using coords (3,2)
2 = -3(3) + c
c = 11
y = -3x + 11.
Hope this helps.
:-)